# thanks: Author to whom correspondence should be addressed

### v1

@@ 2.3 Model verification
In this subsection, we verify (i.e., perform a self-consistency check on) the proposed $f_{gl}$ correlations given in Eqs. (6). We use $\epsilon=0.345$ and $d_{p}=3~{}$\mathrm{mm}\text{/}$$ in our calculations to be consistent with the PBRE dataset (i.e., the experiments of Motil et al. (2021)). Specifically, we resolve Eq. (5), which was obtained by eliminating the pressure gradient from the momentum equations (3), by treating $\beta$ (in $f_{gl}$) as one of the inputs (along with $v_{ls}$ and $v_{gs}$) and using Matlab’s fsolve routine (with default settings, such as tolerances) to solve the nonlinear equation (5), which yields $\phi_{l}$ as an output. The process is repeated for each $Re^{*}_{gs}$ and $Re^{*}_{ls}$. Then, $-dp/dz=A_{ls}/\phi_{l}^{2}$ (recall Sec. 2.1), which is compared to the PBRE dataset values of $-dp/dz$ in Fig. 3. As seen from Fig. 3, the pressure drops for both the packing materials match exactly, as expected, because the PBRE $-dp/dz$ data was used to infer $\beta$. This comparison verifies the self-consistency of the approach by which we obtained the $f_{gl}$ correlations. Based on this approach, the next section provides a validation of the correlations.
@@ 3 Example implementation of the proposed $f_{gl}$ in an Euler-Euler CFD simulation
One approach to CFD simulation of a PBR is to explicitly resolve the gas-liquid flow through the tortuous paths created in a three-dimensional (3D) packing (Nagrani, Marconnet, and Christov, 2024; Ambekar, Rüde, and Buwa, 2022a; Ambekar, Mondal, and Buwa, 2021; Ambekar, Rüde, and Buwa, 2022b; Xu et al., 2022). Interface-resolved CFD simulations are computationally expensive, in particular because they require a fine mesh to resolve the interstitial flow domain between the packing structures. Another approach to CFD simulations of PBRs is an Euler-Euler two-fluid formulation (Jakobsen, 2008). In this approach, suitable drag force closures must be provided to a TFM to capture the liquid-solid and liquid-gas interactions. When properly calibrated, Euler-Euler simulations can be used to predict bulk (or integrated) quantities, such as the pressure drop across the PBR. However, the accuracy of this approach requires suitable drag correlations. In this section, we demonstrate that the correlations derived in Sec. 2 from the PBRE data are useful for Euler-Euler CFD formulations.
@@ 3.1 Simulation methodology
We simulate transient two-phase gas-liquid flow through an empty 2D planar channel using an Euler-Euler CFD formulation in ANSYS Fluent 2022 R1 (ANSYS Inc., 2022b). This formulation is based on a TFM. We account for porous medium (and hence porosity) implicitly by incorporating the $f_{gl}$ and $f_{ls}$ closure relations from Sec. 2 using user-defined functions (UDFs).
The 2D channel and the corresponding mesh are shown in Fig. 4. We create a CFD geometry similar to the experimental test facility for NASA’s PBRE(Motil, Mitchell, and Balakotaiah, 2020; Motil et al., 2021). To reduce the computational power required to perform the simulations, however, the 2D channel width is $5.08~{}$\mathrm{cm}\text{/}$$ and its length is $30~{}$\mathrm{cm}\text{/}$$, which is half of the experimental setup’s length (Motil et al., 2021). We believe the flow is fully developed much upstream of this $30~{}$\mathrm{cm}\text{/}$$ region, and our CFD simulations predict linear pressure variation with axial position. Hence, the pressure drop from the CFD simulations can be compared to the data of Motil et al. (2021). In addition, the gas phase is nitrogen, the liquid phase is water, and constant properties are assumed as indicated in Table 1.
@@
The boundary conditions (BCs) imposed on the CFD simulations are shown in Fig. 4. At the inlet, constant mesoscale velocities calculated from the superficial velocities (Gunjal et al., 2005) are specified as $v_{g}(y,z=0,t)=v_{gs}/\phi_{g}(y,z=0,t)$ and $v_{l}(y,z=0,t)=v_{ls}/\phi_{l}(y,z=0,t)$ for all $t\geq 0$ and $y$, assuming $\epsilon=1$ at the inlet plane. Since Motil et al. (2021) do not report the inlet volume fraction for the phases, we assumed perfectly mixed conditions and assigned $\phi_{g}(y,z=0,t)=0.5$ for all $t\geq 0$ and $y$. A zero pressure gradient is assigned at the outlet. No-slip velocity conditions ($v_{g}=v_{l}=0$) and $\partial\phi/\partial y=0$ are imposed at the channel’s top and bottom walls. As mentioned before, the porous medium is accounted for implicitly via the interphase coupling drag terms, $f_{gl}$ from Eq. (6) and $f_{ls}$ from Eq. (4). The source terms for each phase’s axial momentum equation are defined as
$f_{g}$ $=-f_{gl},$ (7a)
$f_{l}$ $=f_{gl}-f_{ls},$ (7b)
and are implemented in Fluent using a UDF based on local velocities (instead of the superficial velocities used in Eq. (4) and Eq. (6)), following previous CFD studies (Atta, Roy, and Nigam, 2007b, a; Lopes and Quinta-Ferreira, 2008; Lu et al., 2018; Gunjal et al., 2005). We assume constant porosity of $\epsilon=0.345$. For initial conditions, we assumed partial liquid holdup and hence $\phi_{g}(y,z,t=0)=0.5$ for all points $(y,z)$ in the domain. We also simulated different initial conditions, for example, $\phi_{l}(y,z,t=0)=0.8$, and found the final CFD pressure drop to be generally independent of the initial liquid holdup $\phi_{l}(y,z,t=0)$.
@@
In ANSYS Fluent (ANSYS Inc., 2022b), we used a “Least Squared Cell Based” numerical scheme to discretize gradients. The pressure equation was discretized using the “PRESTO!” scheme and “First Order Upwind” schemes were used to discretize the momentum and volume fraction equations, while a “First Order Implcit” transient formulation was employed to ensure the stability of the solver. Finally, the pressure-velocity coupling was achieved by the “Phase Coupled SIMPLE” algorithm. To ensure the stability and convergence of the solver, we treated the source terms semi-implicitly, meaning the constant part of the source term is explicit, and the terms involving the velocity are implicit in the momentum equations (see Ref. ANSYS Inc., 2022a for detailed explanation).
@@ 3.2 Comparison of pressure drop in PBRs
We performed CFD simulations for Teflon and glass packings, each across different $Re^{*}_{gs}$ and $Re^{*}_{ls}$ values. Next, we compare the pressure drop from the CFD simulations to the experimental pressure drops in PBRE dataset. We restrict our attention to the laminar regime of low $Re^{*}_{gs}$ and $Re^{*}_{ls}$ because we did not incorporate turbulence modeling in the simulations. However, it must be noted that we use the entire data set (Motil, Mitchell, and Balakotaiah, 2020) to calculate $\beta$ and obtain the relevant $f_{gl}$ correlations (recall Sec. 2). The CFD simulations run until $\phi_{l}$ and $-dp/dz$ stop varying with time, and hence, a steady state has been established.
@@
Figure 5 shows the variation of $-dp/dz$ with $Re^{*}_{gs}$ at different $Re^{*}_{ls}$ for (a) Teflon and (b) glass packing conditions. For the Teflon packing, in Fig. 5(a), the CFD simulation is performed for $Re^{*}_{ls}=21.0$, $56.0$, $70.0$, and $20.1<Re^{*}_{gs}<$ 2. The CFD simulations predict a $-dp/dz$ within $\approx 25\%$ of the experimental values, highlighting the validity of the proposed approach based on the Euler-Euler simulation using the proposed $f_{gl}$ correlation for the Teflon packing. The highest error of $\approx 40\%$ occurs at the lowest flow rates and is an artifact of the small pressure drops under consideration. The PBRE datasets(Motil, Mitchell, and Balakotaiah, 2020) do not report the uncertainty of all experimental measurements.
@@ 4 Conclusion
Motivated by the lack of studies of two-phase gas-liquid flow through a PBR at microgravity conditions, we developed a data-driven methodology for the calibration of a gas-liquid drag, $f_{gl}$, closure that can be used in CFD simulations. Starting from a 1D TFM, we proposed a novel closure relation for $f_{gl}$ for two different packing materials having different wettability characteristics — Teflon (nonwetting) and glass (wetting). Specifically, we assumed fully-developed 1D flow at steady-state with cross-sectionally uniform velocities and shared pressures between the gas and liquid phases, thus rendering the TFM into a single Eq. (5). Using the latter, we calculated a drag coefficient $\beta$, such that $f_{gl}=\beta(v_{g}-v_{l})$, for each value of the gas and liquid Reynolds numbers, $Re^{*}_{gs}$ and $Re^{*}_{ls}$, available in the PBRE dataset from NASA’s PSI system (Motil, Mitchell, and Balakotaiah, 2020). Then, we performed composite curve fitting to deduce correlations for $f_{gl}$ for Teflon and glass packing, given in Eqs. (6). To the best of our knowledge, the proposed $f_{gl}$ correlations are one of the first data-driven closure relations to be developed for gas-liquid flows through a PBR at microgravity conditions.

### v2

@@ 2.3 Model verification
In this subsection, we verify (i.e., perform a self-consistency check on) the proposed $f_{gl}$ correlations given in Eqs. (6). We use $\epsilon=0.345$ and $d_{p}=3~{}$\mathrm{mm}\text{/}$$ in our calculations to be consistent with the PBRE dataset (i.e., the experiments of Motil et al. 9). Specifically, we resolve Eq. (5), which was obtained by eliminating the pressure gradient from the momentum equations (3), by treating $\beta$ (in $f_{gl}$) as one of the inputs (along with $v_{ls}$ and $v_{gs}$) and using Matlab’s fsolve routine (with default settings, such as tolerances) to solve the nonlinear equation (5), which yields $\phi_{l}$ as an output. The process is repeated for each $Re^{*}_{gs}$ and $Re^{*}_{ls}$. Then, $-dp/dz=A_{ls}/\phi_{l}^{2}$ (recall Sec. 2.1), which is compared to the PBRE dataset values of $-dp/dz$ in Fig. 3. As seen from Fig. 3, the pressure drops for both the packing materials match well, as expected, because the PBRE $-dp/dz$ data was used to infer $\beta$. Nevertheless, we observe a small number of outliers at for large values of $Re_{ls}^{*}$ and low values of $Re_{gs}^{*}$. We hypothesize the mismatch for these few points is because the assumptions undertaken to develop the model, such as steady state and fully developed flow, may not hold in this regime with potentially unstable or unsteady flow existing, thus even though the value of $\beta$ (and hence $f_{gl}$) was fit to the data, the 1D TFM model does not reproduce it. In the present study, we utilize the proposed correlations in the limit of low $Re_{ls}^{*}$ and low $Re_{gs}^{*}$ where the model predictions are accurate. This comparison verifies the self-consistency of the approach by which we obtained the $f_{gl}$ correlations. Based on this approach, the next section provides a validation of the correlations.
@@ 3 Example implementation of the proposed $f_{gl}$ in an Euler-Euler CFD simulation
One approach to CFD simulation of a PBR is to explicitly resolve the gas-liquid flow through the tortuous paths created in a three-dimensional (3D) packing 24, 28, 29, 30, 31. Interface-resolved CFD simulations are computationally expensive, in particular because they require a fine mesh to resolve the interstitial flow domain between the packing structures. Another approach to CFD simulations of PBRs is an Euler-Euler two-fluid formulation 32. In this approach, suitable drag force closures must be provided to a TFM to capture the liquid-solid and liquid-gas interactions. When properly calibrated, Euler-Euler simulations can be used to predict bulk (or integrated) quantities, such as the pressure drop across the PBR, and not interstitial dynamics or the flow regimes. However, the accuracy of this approach requires suitable drag correlations. In this section, we demonstrate that the correlations derived in Sec. 2 from the PBRE data are useful for Euler-Euler CFD formulations.
@@ 3.1 Simulation methodology
We simulate transient two-phase gas-liquid flow through an empty 2D planar channel using an Euler-Euler CFD formulation in ANSYS Fluent 2022 R1 33. This formulation is based on a TFM. We account for porous medium (and hence porosity) implicitly by incorporating the $f_{gl}$ and $f_{ls}$ closure relations from Sec. 2 using user-defined functions (UDFs).
In the context of a TFM, which by definition does not resolve the details of the interstitial gas-liquid dynamics, the flow in a 3D PBR would be treated as axisymmetric, with suitable drag terms to take into account the liquid-solid and liquid-gas interactions. A 3D axisymmetric flow is essentially 2D, with an axial and cross-section coordinate. Thus, for simplicity but without loss of generality, we simulate a 2D channel geometry to test our approach to fitting the drag terms. The 2D channel and the corresponding mesh are shown in Fig. 4. We create a CFD geometry similar to the experimental test facility for NASA’s PBRE25, 9. To reduce the computational power required to perform the simulations, however, the 2D channel width is $5.08~{}$\mathrm{cm}\text{/}$$ and its length is $30~{}$\mathrm{cm}\text{/}$$, which is half of the experimental setup’s length 9. We believe the flow is fully developed much upstream of this $30~{}$\mathrm{cm}\text{/}$$ region, and our CFD simulations predict linear pressure variation with axial position. Hence, the pressure drop from the CFD simulations can be compared to the data of Motil et al. 9. In addition, the gas phase is nitrogen, the liquid phase is water, and constant properties are assumed as indicated in Table 1.
@@
The boundary conditions (BCs) imposed on the CFD simulations are shown in Fig. 4. At the inlet, constant mesoscale velocities calculated from the superficial velocities 23 are specified as $v_{g}(y,z=0,t)=v_{gs}/\phi_{g}(y,z=0,t)$ and $v_{l}(y,z=0,t)=v_{ls}/\phi_{l}(y,z=0,t)$ for all $t\geq 0$ and $y$, assuming $\epsilon=1$ at the inlet plane. Since Motil et al. 9 do not report the inlet volume fraction for the phases, we assumed perfectly mixed conditions and assigned $\phi_{g}(y,z=0,t)=0.5$ for all $t\geq 0$ and $y$. A zero pressure gradient is assigned at the outlet. No-slip velocity conditions ($v_{g}=v_{l}=0$) and $\partial\phi/\partial y=0$ are imposed at the channel’s top and bottom walls. As mentioned before, the porous medium is accounted for implicitly via the interphase coupling drag terms, $f_{gl}$ from Eq. (6) and $f_{ls}$ from Eq. (4). The source terms for each phase’s axial momentum equation are defined as
$f_{g}$ $=-f_{gl},$ (7a)
$f_{l}$ $=f_{gl}-f_{ls},$ (7b)
and are implemented in ANSYS Fluent using a UDF based on local velocities (instead of the superficial velocities used in Eq. (4) and Eq. (6)), following previous CFD studies 19, 20, 21, 22, 23. We assume constant porosity of $\epsilon=0.345$. For initial conditions, we assumed partial liquid holdup and hence $\phi_{g}(y,z,t=0)=0.5$ for all points $(y,z)$ in the domain. We also simulated different initial conditions, for example, $\phi_{l}(y,z,t=0)=0.8$, and found the final CFD pressure drop to be generally independent of the initial liquid holdup $\phi_{l}(y,z,t=0)$.
In ANSYS Fluent 33, we used a “Least Squared Cell Based” numerical scheme to discretize gradients. The pressure equation was discretized using the “PRESTO!” scheme and “First Order Upwind” schemes were used to discretize the momentum and volume fraction equations, while a “First Order Implcit” transient formulation was employed to ensure the stability of the solver. Finally, the pressure-velocity coupling was achieved by the “Phase Coupled SIMPLE” algorithm. To ensure the stability and convergence of the solver, we treated the source terms semi-implicitly, meaning the constant part of the source term is explicit, and the terms involving the velocity are implicit in the momentum equations (see Ref. 34 for detailed explanation).
@@ 3.2 Comparison of pressure drop in PBRs
We performed CFD simulations for Teflon and glass packings, each across different $Re^{*}_{gs}$ and $Re^{*}_{ls}$ values. Next, we compare the pressure drop from the CFD simulations to the experimental pressure drops in PBRE dataset. We restrict our attention to the laminar regime of low $Re^{*}_{gs}$ and $Re^{*}_{ls}$ because we did not incorporate turbulence modeling in the simulations. However, it must be noted that we use the entire data set 25 to calculate $\beta$ and obtain the relevant $f_{gl}$ correlations (recall Sec. 2). The CFD simulations run until $\phi_{l}$ and $-dp/dz$ stop varying with time, and hence, a steady state has been established.
@@
Figure 5 shows the variation of $-dp/dz$ with $Re^{*}_{gs}$ at different $Re^{*}_{ls}$ for (a) Teflon and (b) glass packing conditions. For the Teflon packing, in Fig. 5(a), the CFD simulation is performed for $Re^{*}_{ls}=21.0$, $56.0$, $70.0$, and $20.1<Re^{*}_{gs}<$ 2. The CFD simulations predict a $-dp/dz$ within $\approx 25\%$ of the experimental values, highlighting the validity of the proposed approach based on the Euler-Euler simulation using the proposed $f_{gl}$ correlation for the Teflon packing. The highest error of $\approx 40\%$ occurs at the lowest flow rates and is an artifact of the small pressure drops under consideration. The PBRE datasets25 do not report the uncertainty of all experimental measurements.
@@ 4 Conclusion
Motivated by the lack of studies of two-phase gas-liquid flow through a PBR at microgravity conditions, we used PBRE data from NASA’s PSI system 25, which accesses only the bubble and pulse flow regimes at microgravity for monodisperse Teflon (nonwetting) and glass (wetting) packing materials, to develop a closure relation for $f_{gl}$ for two different packing materials. Based on the predominant approach in the literature, we assumed fully-developed 1D flow at steady-state with cross-sectionally uniform velocities and shared pressure gradients between the gas and liquid phases, thus rendering the TFM into a single Eq. (5). Using the latter, we followed a data-driven methodology to calculate a drag coefficient $\beta$, featured in the gas-liquid drag, $f_{gl}=\beta(v_{g}-v_{l})$, for each value of the gas and liquid Reynolds numbers, $Re^{*}_{gs}$ and $Re^{*}_{ls}$, available in the PBRE dataset. Then, we performed composite curve fitting to form the correlations for $f_{gl}$ for Teflon and glass packing, given in Eqs. (6).
@@
{acknowledgement}
{suppinfo}
The data supporting this study’s findings are openly available in the Purdue University Research Repository (PURR) at http://doi.org/10.4231/S34E-8N27.
The zip archive named “Math_model_figs.zip” contains the MATLAB scripts for implementing the model described in Sec. 2, and the code to make Figs. 1, 2, and 3. This zip archive also contains the pressure drop data downloaded from NASA’s PSI system and the post-processed CFD data from the present study. The code to generate Fig. 5 is also included.
The zip archives named “Glass_packing.zip” and “Teflon_packing.zip” contain the CFD simulation case files along with the UDF for ANSYS Fluent, as described in Sec. 3.